# SemShareKV: Efficient KVCache Sharing for Semantically Similar Prompts via Token-Level LSH Matching

### v1

@@ 5.2 Benchmarking Evaluation
We argue that using Fuzzy Token Match introduces only a negligible overhead to model inference. Table 1 reports the ROUGE-L scores (Lin, 2004). Benchmarking results show that SemShareKV achieves performance comparable to or better than other baseline methods. Notably, in 4 out of the 5 evaluated datasets, Fully Recompute fails to attain the highest performance scores. We attribute this phenomenon to the token eviction mechanisms employed by SemShareKV, SnapKV, and H2O. By selectively retaining only the most semantically significant tokens for self-attention computation, these methods effectively reduce redundant information in the semantic representation, thereby enhancing the model’s generation quality.
@@
Figure 10: Impact of Prompt Similarity
Figure 11: The three retention patterns start from the same retention ratio.
@@ 5.5 Ablation Study
We conducted two ablation studies to evaluate the impact of fuzzy token matching on semantic understanding. First, by either zeroing out or replacing matched KV cache tokens with random ones, we observed significantly lower ROUGE-L scores compared to the full SemShareKV method, confirming the importance of fuzzy matching for capturing semantics (Table 2). Second, analyzing the KVCache compression ratio on the MultiNews dataset (Figure 9), retaining too much cache adds redundancy and harms performance, while retaining too little causes information loss, emphasizing the need for a balanced cache retention strategy.
@@ 6 Conclusion
While SemShareKV effectively shares KV caches for semantically similar prompts, its evaluation so far has been limited to summarization tasks. Speedups decrease for shorter prompts due to the overhead of fuzzy matching, and several hyperparameters require careful tuning. Additionally, our current implementation focuses on demonstrating SemShareKV’s effectiveness and does not yet support FlashAttention; we plan to incorporate it in future work. Also, the matching threshold is empirically set, and exploring adaptive strategies remains future work.
@@ Appendix A Formula and Inference
Table B1: Similarity evaluation on benchmarking datasets using ROUGE-L, BERTScore, and BLEU. All datasets contain 100 semantically similar rewritten samples.
@@ A.2 Locality-Sensitive Hashing (LSH)
Locality-Sensitive Hashing (LSH) is a technique that enables efficient approximate nearest neighbor searches in high-dimensional spaces by ensuring similar input items are hashed into the same bucket with high probability (Indyk and Motwani, 1998). This reduces the number of distance computations required, making LSH particularly useful for large datasets in applications such as image retrieval and natural language processing (Datar et al., 2004). In LSH for Euclidean distance, a common hash function is:
@@
The LSH (Locality-Sensitive Hashing) in SemShareKV is implemented using the FAISS Python library (Douze et al., 2024). Further configuration details can be found in the provided code repository.
@@ A.4 Key-Value Deviation
We define Key-Value Deviation with L2 norm as below:
@@ B.1 Benchmark Datasets
    MultiNews (Bai et al., 2023): This datasets contain samples composed of multiple independent passages or articles. To generate semantically similar samples, we randomly select one passage or article from each sample and use the Llama 3 model to rewrite it while preserving the original semantics. The rewritten passage is constrained to have a similar length to the original (within a 10% difference in token count). We then replace the original passage with the rewritten one to construct a semantically similar prompt. The position of the rewritten passage naturally varies across samples, appearing at the beginning, middle, or end of the context.
@@
    SAMSum (Gliwa et al., 2019), PubMed, BigPatent, BookSum (Kwan et al., 2023), LCC (Guo et al., 2023): These datasets consist of semantically continuous text or codes. For each sample, we divide the context into individual sentences and randomly select a contiguous segment of the total sentence count. This segment is rewritten using the Llama 3 model, with the constraint that the token count deviates by less than 10% from the original. The rewritten segment replaces the original to create a semantically similar prompt, with its position varying within the context in a similar manner.
@@
    Qasper (Dasigi et al., 2021) and WikiHow (Koupaee and Wang, 2018): These datasets consist of Q&A tasks where each question must be answered based on a specific provided context. To preserve the accuracy of the questions, we use the LLM to rewrite only part of the context, leaving the questions unchanged.
@@
    MMLU (Hendrycks et al., 2021a): MMLU is a multiple-choice question-answering dataset. To ensure the logical integrity of the questions and preserve the original answers, we prompt the LLM to paraphrase each entire question.
Table B1 presents the results of the similarity evaluation, measured using ROUGE-L (Lin, 2004), BLEU (Papineni et al., 2002), and BERTScore (Zhang et al., 2019). We include both longest common subsequence-based metrics (ROUGE-L), n-gram-based metrics (BLEU) and embedding-based metrics (BERTScore) to provide a comprehensive evaluation of semantic similarity across rewritten datasets.
@@ Appendix C Extra Experimental Results
Figure C2 shows the three retention patterns mentioned in Insights 3.
@@ Appendix D Implementation and Hyperparameters
SemShareKV is implemented in Python using the transformers library (Wolf et al., 2020), with the monkeypatching technique. We use the Locality Sensitive Hashing from FAISS Douze et al. (2024) library. The code is available at: https://anonymous.4open.science/r/SemShareKV-B53C. Details of the key functions and their roles are outlined below:
@@
Overall, SemShareKV is built on the transformer architecture and consists of fewer than 300 new lines of code, making it lightweight and easily transferable to other LLMs.

### v2

@@ 5.2 Benchmarking Evaluation
We argue that using Fuzzy Token Match introduces only a negligible overhead to model inference. Table 1 reports the ROUGE-L scores (lin2004rouge). Benchmarking results show that SemShareKV achieves performance comparable to or better than other baseline methods. Notably, in most of the evaluated datasets, Fully Recompute fails to attain the highest performance scores. We attribute this phenomenon to the token eviction mechanisms employed by SemShareKV, SnapKV, PyramidKV and H2O. By selectively retaining only the most semantically significant tokens for self-attention computation, these methods effectively reduce redundant information in the semantic representation, thereby enhancing the model’s generation quality.
@@ 5.5 Ablation Study
We conducted three ablation studies to assess the effect of fuzzy token matching on semantic understanding. First, when applying the fuzzy matching technique with the full cache, we observed a decline in performance, indicating the necessity of the token retention mechanism. Second, when matched KV cache tokens were either zeroed out or replaced with random ones, the ROUGE-L scores dropped significantly compared to the full SemShareKV method, confirming that fuzzy matching is crucial to capture semantic relationships (Table 2). Third, as shown in Figure 10, an analysis of the KV cache compression ratio in the MultiNews dataset revealed that retaining too much cache introduces redundancy and degrades performance, while retaining too little leads to information loss, highlighting the importance of a balanced cache retention strategy.
@@ 6 Conclusion
Although SemShareKV effectively shares KV caches across semantically similar prompts, the speedup decreases for shorter prompts because of the overhead introduced by fuzzy token matching, and several hyper-parameter require careful tuning. Our current implementation focuses on demonstrating the effectiveness of SemShareKV and does not yet support FlashAttention, which we plan to explore into future work. The matching threshold is also empirically determined, and exploring adaptive strategies remains an open direction. In the ablation study, we replaced sentences with other samples from the same dataset, ensuring a fair comparison within a consistent data distribution. Although cross-corpus replacement could test broader robustness, our setup effectively isolates the impact of semantic mismatches. This opens future work not only to extend the experiments to cross-corpus settings but also to construct datasets of semantically similar prompts to better evaluate robustness.
@@ Appendix A Formula and Inference
Table B1: Similarity evaluation on benchmarking datasets using ROUGE-L, BERTScore, and BLEU.
@@ A.2 Locality-Sensitive Hashing (LSH)
Locality-Sensitive Hashing (LSH) is a technique that enables efficient approximate nearest neighbor searches in high-dimensional spaces by ensuring similar input items are hashed into the same bucket with high probability (indyk1998approximate). This reduces the number of distance computations required, making LSH particularly useful for large datasets in applications such as image retrieval and natural language processing (datar2004locality). In LSH for Euclidean distance, a common hash function is:
@@ A.4 Key-Value Deviation
We define Key-Value Deviation with $L_{2}$ norm as below:
@@ B.1 Benchmark Datasets
    MultiNews (bai2023longbench): This datasets contain samples composed of multiple independent passages or articles. To generate semantically similar samples, we randomly select one passage or article from each sample and use the Llama 3 model to rewrite it while preserving the original semantics. The rewritten passage is constrained to have a similar length to the original (within a 10% difference in token count). We then replace the original passage with the rewritten one to construct a semantically similar prompt. The position of the rewritten passage naturally varies across samples, appearing at the beginning, middle, or end of the context.
@@
    SAMSum (gliwa-etal-2019-samsum), PubMed, BigPatent, BookSum (kwan_m4le_2023), LCC (guo2023longcoder): These datasets consist of semantically continuous text or codes. For each sample, we divide the context into individual sentences and randomly select a contiguous segment of the total sentence count. This segment is rewritten using the Llama 3 model, with the constraint that the token count deviates by less than 10% from the original. The rewritten segment replaces the original to create a semantically similar prompt, with its position varying within the context in a similar manner.
@@
    Qasper (Dasigi2021ADO) and WikiHow (koupaee2018wikihow): These datasets consist of Q&A tasks where each question must be answered based on a specific provided context. To preserve the accuracy of the questions, we use the LLM to rewrite only part of the context, leaving the questions unchanged.
@@
    MMLU (hendrycks2021ethics): MMLU is a multiple-choice question-answering dataset. To ensure the logical integrity of the questions and preserve the original answers, we prompt the LLM to paraphrase each entire question.
Table B1 presents the results of the similarity evaluation, measured using ROUGE-L (lin2004rouge), BLEU (papineni2002bleu), and BERTScore (zhang2019bertscore). We include both longest common subsequence-based metrics (ROUGE-L), n-gram-based metrics (BLEU) and embedding-based metrics (BERTScore) to provide a comprehensive evaluation of semantic similarity across rewritten datasets.
@@ Appendix C Extra Experimental Results
Figure C2 illustrates the three retention patterns discussed in Insights 3. Table B2 provides additional benchmarking results of Total Batch Time (TBT) and throughput. Since all experiments were conducted with batch size equal to 1, the reported throughput values are equivalent to TBT.
Table B2: Effieicny Comparision with Throughput
Method Input Length Token / Sec)
Fully Recompute 5k 12.681
SemShareKV 5k 34.065
SnapKV 5k 22.576
H2O 5k 16.273
PyramidKV 5k 3.991
@@ Appendix D Implementation and Hyperparameters
SemShareKV is implemented in Python using the transformers library (wolf-etal-2020-transformers), with the monkeypatching technique. We use the Locality-Sensitive Hashing from FAISS douze2024faiss library. The code is available at: https://github.com/JasperZhao666/SemShareKV-public.git. Details of the key functions and their roles are outlined below:
@@
In general, SemShareKV is built on the transformer architecture and consists of fewer than 300 new lines of code, which makes it lightweight and easily transferable to other LLMs.